Kate has n keys, only one of which will open the cookie jar. (a) If she tries the keys at random, discarding those that do not work, what is the (b) Later in the day, her brother Kurtis wants to get a cookie, but he does not probability that she will open the jar on her kth try? discard keys that he has previously tried. What is the probability that he will open the jar on his kth try? (c) Kurtis forgets to re-lock it, and his sister Caroline finds the open cookie jar Caroline knows that the time until the next person walks into the kitchen follows an exponential distribution with rate A-.95. It takes her 2 minutes to eat the cookie. What is the probability that someone catches her sneaking a cookie? (d) Given that no one has walked in during the two minutes Caroline was eating her first cookie, what is the probability that someone will catch her while she eats her second cookie?
Answer #1

(a)

Probability that she will get the correct keys on xth try = (n - x) / (n-x + 1) (Because n-x+1 keys are available and n-x keys are not the correct keys)

Probability that she will open the jar on kth try = Probability that she will get incorrect keys on (k-1)'s tries * Probability that she will get correct keys on kth try

= (n-1/n) * (n-2)/(n-1) * ... * (n - k + 1) / (n-k+2) * (1 / (n - k + 1)

= 1 / n

(b)

If Kurtis do not discard keys, the probability that he will get correct keys on any try = 1/n

Probability that he will open the jar on kth try = Probability that he will get incorrect keys on (k-1)'s tries * Probability that he will get correct keys on kth try

= [(n - 1)/n]k-1 * (1/n)

= (n - 1)k-1 / nk

(c)

Let T be the time interval between person arrival into the kitchen. Then T ~ Exp(\lambda = 0.95)

Probability that someone catches her sneaking a cookie = Probability that someone arrives in kitchen in less than 2 minutes

= P(T \le 2) = 1 - exp(-0.95 * 2) = 0.8504

(d)

Due to memoryless property of exponential distribution, probability that someone catches her eating cookie = Probability that someone arrives in kitchen before next 2 minutes

= P(T \le 2) = 1 - exp(-0.95 * 2) = 0.8504